Prove that an open interval in R is an open set and a closed intervals is a closed set
1
Expert's answer
2020-05-19T20:08:13-0400

A subset "E \sube \R" is called open set if every point of "E" is an interior point of "E."

Let "a,b\in \R" be any two real number .

Let "I=(a,b)" be an open interval.

Claim: "I" is open .

Let "y\in I" be an arbitrary real number then "a<y<b" .

Let "c,d" be two real number such that "a<c<y<d<b"

"\therefore (c,d)\sube(a,b)."

Let "r=min(|c-y|,|y-d|)."

Clearly "r>0" as "y\neq c,d."

Define , "r_0=\frac{r}{2}" .

"\therefore N_{r_0}(y)=\{ x:|x-y|<r_0\}" "=-r_0<x-y<r_0"

"=y-r_0<x<y+r_0=( y-r_0,y+r_0)"

"=\text{Neighborhood of y }"

Let "x\in N_{r_0}(y)" then "y-r_0<x<r_0+y" .

But "r_0<r" "\implies c<y-r_0<y \ and \ y<y+r_0<d" Therefore , "c<x<d."

Hence , "x\in (c,d)."

"\therefore x\in N_{r_0}(y)\sube (c,d)\sube (a,b)."

Therefore every point of "(a,b)" is an interior point of "(a,b)."

Hence , "(a,b)" is open.

Consider the close interval "E=[a,b]" ,where "a,b\in \R" .

We known that a subset "E" of "\R" is close iff "E^c" is open.

Now , "E^c=\{x:x\notin [a,b]\}" "=\{ x: x>b \ or \ x<a\}"

"=(-\infin,a)\cup(b,\infin)"

Clearly "(-\infin,a) \ and \ (b,\infin)" are open as they are open interval.

Again we known that Union of two open set is open ,hence "E^c" is open .

Therefore , "E" is close.